OpenStudy (thecatman):

Parker earned $26 by selling 20 glasses of fresh-squeezed lemonade at his lemonade stand on a Saturday morning. He charged $1.00 for a small glass and $1.50 for a large glass. Use a model to determine how many large glasses he sold. 8 10 12 14

OpenStudy (thecatman):

i think c

OpenStudy (paki):

hm... Let x be the number of large glasses. Since what the total number of glasses used is 20. The small glasses is 20 - x The model equation is: 26 = 1.5(x) + 1(20 - x)... now solve this for "x"... 26 = 1.5x + 20 - 1x 6 = 1.5 - 1x 6 = 0.5x x = 6/(0.5) x = 12

OpenStudy (thecatman):

thanks for helping me

OpenStudy (paki):

always pleasure...